Two capacitors give an equivalent capacitance of 9.02 pF when connected in parallel and an equivalent capacitance of 1.78 pF when connected in series. What is the capacitance of each capacitor? smaller capacitance pF larger capacitance pF
Answer #1

when connected in parallel,

C1 + C2 = 9.2 pF

C1 = 9.2 - C2

when connected in series,

1 / C1 + 1 / C2 = 1 / 1.78

1 / (9.2 - C2) + 1 / C2 = 1 / 1.78

C2^2 - 9.2*C2 + 9.2*1.78 = 0

by solving the quadratic equation,

C2 = 2.41 pF (smaller)

C1 = 9.2 - 2.41 = 6.79 pF (larger)